Let V represent the volume of a cylinder, h represent its height, and r represent its radius. V, h, and r are related according to the formula
r=√v/πh

a)A cylindrical soup can has a volume of 281rin3 It is 7in high. What is the radius of the can?

b)Solve the equation for V.



Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52778)

.

Let me answer question a) only.

My answer is THIS :


    Substitute the given data into the formula and calculate.

    If it is not clear in full to you, let me add: calculate it ON YOUR OWN.


Making calculations is students' honor duty and a privilege !





Answer by Alan3354(69443)

Let V represent the volume of a cylinder, h represent its height, and r represent its radius. V, h, and r are related according to the formula
r=√v/πh

a)A cylindrical soup can has a volume of 281rin3 It is 7in high. What is the radius of the can?
____
h = 7
Vol = pi*r^2*h
Solve for r
=================

b)Solve the equation for V.
V is given.
V = 281*r (maybe????)
sloppy posting.